OpenStudy (anonymous):

Which is true about the graphs of the two functions f(x) = (x-4)(x-9)and f(x)=(x-6)^2

OpenStudy (anonymous):

They have the same range. They have the same vertex. They have the same y-intercept. They have the same axis of symmetry

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Hmmmm

OpenStudy (anonymous):

The y intercept is the easiest to check. Just let \(x=0\).

OpenStudy (anonymous):

What Does That Mean

OpenStudy (anonymous):

\[ (x-4)(x-9)\bigg|^{x=0} = (0-4)(0-9) = -4\times -9 = 36 \]and \[ (x-6)^2\bigg|^{x=0} = (0-6)^2=-6\times -6 = 36 \]

OpenStudy (anonymous):

They both have a \(y\) intercept at \(36\).

OpenStudy (anonymous):

So They Share The Same Y Intercept

OpenStudy (anonymous):

You have to try a bit, dude.

OpenStudy (anonymous):

iAM

OpenStudy (anonymous):

Yeah, they share the same y intercept

OpenStudy (anonymous):

What is the axis of symmetry of the function f(x) =-(x+9)(x-21)